OpenStudy (anonymous):

Help with 2 questions

OpenStudy (anonymous):

simplify the polynomial. 1. 3x^2 + 6 – 2x + 5x – 4x^2 + 9 –x^2 + 3x + 15 7x^2 + 3x + 3 x^2 – 3x + 15 –x^2 + 7x + 15 2. 2x^2 + 6x – 7x + 8 – 3x^2 + 1 2x^2 + x + 9 –2x^2 – x – 9 –x^2 – x + 9 x^2 + 9

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@satellite73

mathslover (mathslover):

Hint : Combine Like terms.

mathslover (mathslover):

3x^2 + 6 – 2x + 5x – 4x^2 + 9 \(3x^2 - 4x^2 + 6 + 9 -2x + 5x\) Simplify it then,can yo?

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

Can you help @Euler271

OpenStudy (anonymous):

what mathslover is doing is the correct approach. you can add the coefficients of the x's of the same order 2x^2 + 3x -2x^2 - 2x = x

OpenStudy (anonymous):

coefficients being the number in front 3x^2 - 4x^2 = (3 - 4)x^2

OpenStudy (texaschic101):

combine you x^2's....then combine your x's...then combine your numbers.....do not try and combine x^2 with x or a number with x^2.....now you try

OpenStudy (anonymous):

2x+5x=7x

OpenStudy (anonymous):

6+9=15

OpenStudy (anonymous):

combining x^2 -x^2

OpenStudy (anonymous):

so together it would be –x^2 + 7x + 15

OpenStudy (anonymous):

can anybody tell me if this is correct?

OpenStudy (anonymous):

then for the second question

OpenStudy (mertsj):

-2x+5x= what?

OpenStudy (mertsj):

if you have 5 dollars and spend 2 dollars you now have 7 dollars?????

OpenStudy (anonymous):

i mean 3x

OpenStudy (mertsj):

Then rethink your answer for number 1

OpenStudy (anonymous):

–x^2 + 3x + 15

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

i don't get the second one

OpenStudy (anonymous):

I know it is not B

OpenStudy (mertsj):

It is the same as the first one....add the like terms

OpenStudy (anonymous):

–x^2 – x + 9?

OpenStudy (mertsj):

yes. good job

OpenStudy (anonymous):

thx!!

OpenStudy (mertsj):

yw